# Causal Interventions in Bond Multi-Dealer-to-Client Platforms

### v1

@@ Abstract
The digitalization of financial markets has shifted trading from voice to electronic channels, with Multi-Dealer-to-Client (MD2C) platforms now enabling clients to request quotes (RfQs) for financial instruments like bonds from multiple dealers simultaneously. In this competitive landscape, dealers cannot see each other’s prices, making a rigorous analysis of the negotiation process crucial to ensure their profitability. This article introduces a novel general framework for analyzing the RfQ process using probabilistic graphical models and causal inference. Within this framework, we explore different inferential questions that are relevant for dealers participating in MD2C platforms, such as the computation of optimal prices, estimating potential revenues and the identification of clients that might be interested in trading the dealer’s axes. We then move into analyzing two different approaches for model specification: a generative model built on the work of (Fermanian, Guéant, & Pu, 2017); and discriminative models utilizing machine learning techniques. We evaluate these methodologies using predictive metrics designed to assess their effectiveness in the context of optimal pricing, highlighting the relative benefits of using models that take into account the internal mechanisms of the negotiation process.
@@ 1 Introduction
Modern financial markets are increasingly mediated by electronic platforms that automate the negotiation and execution of trades. While highly liquid instruments such as equities are typically traded on order-driven venues using continuous double auctions and limit order books, these mechanisms are less effective for less frequently traded instruments, such as corporate and government bonds. Unlike equities, where a company issues a single line of stock, bond issuers may maintain hundreds of distinct securities with varying maturities, coupons, and seniorities. This proliferation of instruments fragments liquidity and reduces the probability of matching counterparties through centralized order books.
@@
This paper introduces a unified framework based on probabilistic graphical models [2, 3, 4] and causal inference [5, 6] to formalize the RfQ process in MD2C platforms and address these challenges. Graphical models allow us to explicitly encode the dependencies among variables involved in the RfQ workflow, including observed and latent factors. Causal inference techniques then enable the analysis of interventions—such as setting specific quotes or initiating outreach to clients—in a manner that accounts for confounding biases and supports counterfactual reasoning. Within this framework, we focus on three high-impact dealer use cases: (i) optimal pricing to maximize expected profits or satisfy commercial targets, (ii) revenue potential estimation under fixed pricing policies, and (iii) identifying likely matches between axes and client demand.
We develop two modeling strategies to instantiate the graphical model: a generative model and a discriminative alternative. The generative model, based on the work by Fermanian, Guéant, and Pu [7], directly models the internal mechanism of the RfQ negotiation, which allows it to automatically satisfy expected economic constraints like price monotonicity —i.e., that clients prefer more favorable prices, all else being equal, and incorporates in a natural way post-trade information, such as the cover price. In contrast, discriminative models —such as logistic regression and gradient-boosted decision trees— [8] focus on directly estimating outcome probabilities given observed RfQ features. These models are more flexible and easier to scale, but may struggle to encode such economic constraints or capture effectively post-trade information.
@@ 2 A causal graphical model for the RfQ process
We model the full RfQ process as a graphical model depicted in Figure 1. This model captures key components of the RfQ process, making it well-suited for pricing analysis as well as other commercial applications, such as client targeting for axe recommendations and estimating a dealer’s expected profitability following a transaction.
Fig 1: Causal graphical model for the RfQ process. Shaded nodes correspond to observable variables, whereas blank ones are latent variables.
@@
  • Client information asymmetry (IA): If the client has relevant information about the short-term directionality of the market [9, 10]. There are different reasons why a client could have information asymmetry. From a causal graph structure, we differentiate between two: 1) the client trading activity represents a potential source of market impact, but the RfQ submitted to the dealer often reveals only a subset of the total order. In this case, the information asymmetry impacts the market, which we model as a price drift ($\mu$) variable, and the direct causal effect: $IA\rightarrow\mu$, 2) the client has insider information or advanced analytics models that are relevant for predicting market directionality, again modeled as a price drift variable. In this case, though, there is not a direct link between IA and drift, since the client uses an estimation of the drift, but does not know the actual drift after trading. Since the first case is more general, we will consider it in the rest of the paper, keeping in mind that extra simplifications of the graphical model can be achieved in the second situation.
@@
  • The half-spread ($\delta$) quoted by the dealer. Although dealers quote prices to clients, we will make the hypothesis that client decisions on trading are driven by the spread with respect to a reference mid-price $P_{m}$, for example a platform’s composite price: $P=P_{m}+s\delta$, where $s=1$ when the dealer sells (ask) and $s=-1$ when the dealer buys (bid).
@@
  • End of day flow value ($R_{\text{eod}}$): Since round-trips of buying and selling might involve large time scales for illiquid bonds, dealers also tend to compute revenues at a given time-horizon, e.g. end of day, by using market mid-prices to value inventories. In this case, hence, the revenue is computed as $R_{\text{eod}}=v\delta_{t}+sv(P_{m,T}-P_{m,t})$, where $T$ is the end of day time. Liquidity penalties can sometimes be included in this formulation to penalize unrealized mark-to-market profits, reflecting the potential cost of carrying illiquid assets.
@@ 3 Causal interventions and predictions in the graphical model
which is an implicit formula for the optimal spread for a general functional form for $f(\delta)$. A central insight from Causal Inference theory [5] is that unbiased estimation of the hit probability model critically depends on selecting an appropriate conditioning set. In many cases, a minimal subset of variables can be identified that ensures the validity of traditional probabilistic estimation methods while preserving causal interpretability. Importantly, including all available features indiscriminately can introduce spurious correlations, distorting the estimation of causal effects.
@@
where the hit probability model naturally shows up again. Their estimation, as before, requires conditioning on at least $\sigma$, $BF$, $CF$, and $RF$. In this computation, though, the utility of revenues on hit is still a random variable which depends, according to our model, on the evolution of the mid-price. We model such dependence using a Brownian Motion model whose drift is contingent on the information asymmetry of the client requesting the RfQ. Such conditioning is also necessary in order to evaluate what we will refer to as revenues on hit model with historical data: as it will be detailed in the next section, to satisfy the back-door criterion and remove the do-operator from the expected value, the conditioning set must include at least the volatility $\sigma$, the RfQ features $RF$, and either the latent variable $IA$ or the drift $\mu$. Being latent variables, this necessarily means integrating over their distribution.
For simplicity, we choose ${\mathcal{Z}_{t}}$ to include $BF$ and $CF$ as well, since the extra variables don’t open new back-door spurious paths for the revenues on hit, allowing us to use a single information set for both the hit probability model and the revenues on hit model. We can now proceed to evaluate the model. Let us first assume that the dealer does not expect informational advantage in this RfQ, and there is no other expectation of short-term directionality of the market. We can then use $P_{m,T}-P_{m,t}\sim N(0,\sigma(T-t))$ and compute the optimal spread:
@@
We can extend this model to the case in which the dealer also has a view on the informational advantage of the client. A simple model that captures the influence of the information asymmetry into the market dynamics is: $\hat{\mu}1_{\text{IA}=1}$. The conditioning set to compute revenues in this case does not change, as discussed in the next section. Working out again the utility maximization problem, the optimal spread has an extra additive correction that compensates for the information asymmetry risk:
@@
[…]
@@
Factoring the axe matcher into these two probabilities provides a more tractable formulation. Since, in our model, there is no direct causal effect between the client’s demand (captured by the RfQ variable) and the spread $\delta$, assuming the existence of a conditioning set $\mathcal{Z}_{t}$ that blocks all the spurious correlations between these variables —without inducing new ones— allows us to simplify the first probability. A similar argument holds for the RfQ status ($RS$) with respect to the commercial action variable $call$ in the second probability. Under these assumptions, the overall axe matcher expression can be rewritten as:
@@
$P(\text{hit}|\text{do}(\delta),\text{RfQ},CF,BF,\mathcal{Z}_{t})\times$
@@
and the ACE reads:
$ACE=P(\text{hit}|\text{do}(\delta),\text{RfQ},CF,BF,\mathcal{Z}_{t})\times$
$\left(P(\text{RfQ}|\text{do}(\text{call}=1),\text{axe}=1,CF,BF,\mathcal{Z}_{t})-\right.$
@@
The first term can be directly linked to the hit probability model, since conditioning on RfQ allows us to remove any influence from the variable $axe$: at this point, the client’s decision to trade is assumed to depend solely on the pricing, regardless of whether the dealer is axed or not. To estimate this model using historical data, as previously discussed, we must include the volatility $\sigma$ and the RfQ features $RF$—such as the requested volume and the number of competing dealers $n$—in the conditioning set. However, since these variables are not known in advance in the context of the axe matcher problem, we must integrate over their conditional probability distribution to properly account for their effect.
@@ 4 Causal analysis of the interventional probabilities
The hit probability model, as discussed in the previous section, is central to the business questions examined in this work. Our objective is to evaluate the causal quantity $P(RS=\text{hit}\mid\text{do}(\delta),\text{RfQ},\mathcal{Z}_{t})$. Specifically, we seek to determine the minimal set of variables in $\mathcal{Z}_{t}$ that enables estimation of this expression from historical (i.e., observational) data. According to Causal Inference theory [5], a sufficient—though not necessary—condition is that the set $\mathcal{Z}_{t}$ satisfies the back-door criterion. Given a causal graph and a pair of variables $X$ and $Y$, a set of variables $\mathcal{Z}$ satisfies this criterion if (i) no variable in $\mathcal{Z}$ is a descendant of $X$, and (ii) $\mathcal{Z}$ blocks all back-door paths from $X$ to $Y$—that is, all paths that contain an arrow pointing into $X$. If such a set $\mathcal{Z}_{t}$ exists, we can validly equate the interventional and observational expressions:
@@
  • The paths via the RfQ features $RF$. On the one hand, the path $\delta\leftarrow RF\rightarrow R\leftarrow RS$ is blocked since $R$, the revenues, is a collider with respect to any direct or indirect path connecting it to $RS$. By not conditioning on revenues, these paths are blocked. Paths passing through $BF$, $CF$ and RfQ are blocked due to the conditioning on those variables. Finally we have paths connecting RfQ features to $RS$ via latent variables like $\delta_{\text{res}}$ and $\delta_{\text{dealer}}$, which are open. Again, to block these spurious influences using observable variables we can include $RF$ in the conditioning set.
@@
×$\times P(RF,BF,CF|\text{RfQ})dP(\sigma)$ (24)
@@
$P(R>0|\text{do}(\delta),\text{hit},\text{RfQ},\sigma,RF)=\sum_{\text{IA}=0,1}P(R>0|\delta,\text{hit},IA,\sigma,RF)P(IA|\delta,\text{hit},\text{RfQ},\sigma,RF)$ (25)
@@ 5 Model specification
So far we have use Causal Inference theory to express business causal interventions in terms of probabilities that can be estimated using historical data. When it comes to actually computing those probabilities, we need to provide a model specification. Here, two main routes can be taken:
@@
  • For the conditional distribution of the spreads quoted by other dealers, we adopt the Skew Exponential Power (SEP) distribution, as proposed in [7], applied to the normalized dealer half-spread ($\delta_{\text{dealer}}/\Delta$). To visually assess the appropriateness of this choice, we leverage a structural property of our model: the conditional independence of dealer quotes given their parent features. Under this assumption, the cover price—defined as $\delta_{\text{cover}}=\max\{\delta_{\text{dealer},1},\dots,\delta_{\text{dealer},n}\}$ for buy-side RfQs (and the minimum for sell-side RfQs)—follows the distribution of the maximum of $n$ independent random variables, each drawn from the same SEP distribution conditional on parent features. Figure 2 displays this distributional fit using the SEP assumption. Notably, despite our dataset differing from the one used in [7], the results offer further evidence of the robustness of the SEP specification. To model the conditional distribution of dealer spreads with respect to their parent nodes—which include bond, client, and RfQ features, as well as volatility—we use a linear model:
@@
  • For the distribution of the price discovery variable $PD$, conditional on its parent variables $BF$ and $CF$, one could, in principle, use Bernoulli distributions. However, this approach risks an exponential increase in the number of parameters if a large number of features is included. To address this issue while preserving interpretability and tractability, we model $PD$—a binary random variable—using a logistic regression model, namely:

### v2

@@ Abstract
The digitalization of financial markets has shifted trading from voice to electronic channels, with Multi-Dealer-to-Client (MD2C) platforms now enabling clients to request quotes (RfQs) for financial instruments like bonds from multiple dealers simultaneously. In this competitive landscape, dealers cannot see each other’s prices, making a rigorous analysis of the negotiation process crucial to ensure their profitability. This article introduces a novel general framework for analyzing the RfQ process using probabilistic graphical models and causal inference. Within this framework, we explore different inferential questions that are relevant for dealers participating in MD2C platforms, such as the computation of optimal prices, estimating potential revenues and the identification of clients that might be interested in trading the dealer’s axes. We then move into analyzing two different approaches for model specification: a generative model built on the work of (Fermanian, Guéant, & Pu, 2017); and discriminative models utilizing machine learning techniques. Our results show that generative models can match the predictive accuracy of leading discriminative algorithms such as LightGBM (ROC-AUC: 0.742 vs. 0.743) while simultaneously enforcing critical business requirements, notably spread monotonicity.
@@ 1 Introduction
Modern financial markets are increasingly mediated by electronic platforms that automate the negotiation and execution of trades. While highly liquid instruments such as equities are typically traded on order-driven venues using continuous double auctions and limit order books, these mechanisms are less effective for instruments that are not as actively traded, such as corporate and government bonds. Unlike equities, where a company issues a single line of stock, bond issuers may maintain hundreds of distinct securities with varying maturities, coupons, and seniorities. This proliferation of instruments fragments liquidity and reduces the probability of matching counterparties through centralized order books.
@@
To address the first research question of this study —namely, assessing the potential impact of confounding factors in the historical datasets used by dealers for decision-making—this paper introduces a unified framework based on probabilistic graphical models [2, 3, 4] and causal inference [5, 6]. Graphical models allow us to explicitly encode the dependencies among variables involved in the RfQ workflow, including observed and latent factors. Causal inference techniques then enable the analysis of interventions—such as setting specific quotes or initiating outreach to clients—in a manner that accounts for confounding biases and supports counterfactual reasoning. In particular, we focus on three high-impact dealer use cases: (i) optimal pricing to maximize expected profits or satisfy commercial targets, (ii) revenue potential estimation under fixed pricing policies, and (iii) identifying likely matches between axes and client demand.
Our work thus continues a recent trend in which causal inference methods are being popularized across economics and finance, although their use in financial markets—traditionally dominated by correlation-based analyses—remains at an early stage. Early applications in empirical economics have focused on identifying causal effects in macroeconomic and policy contexts. For example, Baiardi and Naghi [7] revisited classical empirical financial economics studies using Double Machine Learning [8], a methodology that integrates causal inference with modern machine learning to estimate structural relationships more reliably. Similarly, [9] have examined the causal impact of interest rate changes on fixed income funds through Double ML estimation, illustrating the growing use of causal ML to evaluate policy transmission mechanisms. More recently, causal methods have been applied directly to financial markets: Oliveira et al. [10] have proposed causality-inspired forecasting models that remain robust under distributional shifts, demonstrating improved performance in volatile investment environments. Extending this line of research, López de Prado and Zoonekynd [11] advocate for a causal investing paradigm grounded in counterfactual reasoning and experimental design, framing causal inference as a foundation for more reliable discovery of investment factors and strategies.
The second primary research question of this study concerns the comparative effectiveness of different modeling paradigms for representing the RfQ negotiation process. Specifically, we aim to assess the relative advantages of generative models that incorporate the internal mechanics of dealer–client interactions versus discriminative models that directly learn predictive mappings from data. To this end, we develop two complementary strategies to instantiate the graphical model.
The generative model, inspired by the work of Fermanian, Guéant, and Pu [12], explicitly represents the negotiation mechanism underlying RfQ interactions. This approach naturally enforces economically consistent behaviors—such as price monotonicity, whereby clients prefer more favorable quotes—and seamlessly integrates post-trade information like the cover price. In contrast, discriminative models, such as logistic regression and gradient-boosted decision trees [13], directly estimate outcome probabilities conditional on observed RfQ features. These models offer greater flexibility and scalability but may struggle to encode economic constraints or effectively leverage post-trade feedback. Comparing these two approaches enables us to evaluate the trade-off between structural interpretability and predictive performance within the causal inference framework.
@@ 2 A causal graphical model for the RfQ process
We model the full RfQ process as a graphical model depicted in Fig 1. This model captures key components of the RfQ process, making it well-suited for pricing analysis as well as other commercial applications, such as client targeting for axe recommendations and estimating a dealer’s expected profitability following a transaction.
Fig 1: Causal graphical model for the RfQ process. Green nodes correspond to observable variables, whereas blue ones are latent variables. A detailed description of each of the variables is provided in the main text
@@
  • Client information asymmetry (IA): If the client has relevant information about the short-term directionality of the market [14, 15]. There are different reasons why a client could have information asymmetry. From a causal graph structure, we differentiate between two: 1) the client trading activity represents a potential source of market impact, but the RfQ submitted to the dealer often reveals only a subset of the total order. In this case, the information asymmetry impacts the market, which we model as a price drift ($\mu$) variable, and the direct causal effect: $IA\rightarrow\mu$, 2) the client has insider information or advanced analytics models that are relevant for predicting market directionality, again modeled as a price drift variable. In this case, though, there is not a direct link between IA and drift, since the client uses an estimation of the drift, but does not know the actual drift after trading. Since the first case is more general, we will consider it in the rest of the paper, keeping in mind that extra simplifications of the graphical model can be achieved in the second situation. Notice that information asymmetry is closely related to adverse selection effects, a central concern in market microstructure. Recent advances, such as [16], provide new evidence and models for understanding toxic flow and its implications for dealer profitability.
@@
  • The half-spread ($\delta$) quoted by the dealer. Although dealers quote prices to clients, we will make the hypothesis that client decisions on trading are driven by the spread with respect to a reference mid-price $P_{m}$, for example a platform’s composite price: $P=P_{m}+s\delta$, where $s=1$ when the dealer sells (ask) and s=1s=-1 when the dealer buys (bid).
@@
  • End of day flow value ($R_{T}$): Since round-trips of buying and selling might involve large time scales for illiquid bonds, dealers also tend to compute revenues at a given time-horizon, e.g. end of day, by using market mid-prices to value inventories. In this case, hence, the revenue is computed as $R_{T}=v\delta_{t}+sv(P_{m,T}-P_{m,t})$, where $T$ is the end of day time. Liquidity penalties can sometimes be included in this formulation to penalize unrealized mark-to-market profits, reflecting the potential cost of carrying illiquid assets.
@@ 3 Causal interventions and predictions in the graphical model
which is an implicit formula for the optimal spread for a general functional form for $f(\delta)$.
A central insight from Causal Inference theory [5] is that unbiased estimation of the hit probability model critically depends on selecting an appropriate conditioning set. In many cases, a minimal subset of variables can be identified that ensures the validity of traditional probabilistic estimation methods while preserving causal interpretability. Importantly, including all available features indiscriminately can introduce spurious correlations, distorting the estimation of causal effects.
@@
where the hit probability model naturally shows up again. Their estimation, as before, requires conditioning on at least $\sigma$, $BF$, CFCF, and $RF$. In this computation, though, the utility of revenues on hit is still a random variable which depends, according to our model, on the evolution of the mid-price. We model such dependence using a Brownian Motion model whose drift is contingent on the information asymmetry of the client requesting the RfQ. Such conditioning is also necessary in order to evaluate what we will refer to as revenues on hit model with historical data: as it will be detailed in the next section, to satisfy the back-door criterion and remove the do-operator from the expected value, the conditioning set must include at least the volatility $\sigma$, the RfQ features $RF$, and either the latent variable $IA$ or the drift $\mu$. Being latent variables, this necessarily means integrating over their distribution.
For simplicity, we choose $\mathcal{Z}_{t}$} to include $BF$ and $CF$ as well, since the extra variables don’t open new back-door spurious paths for the revenues on hit, allowing us to use a single information set for both the hit probability model and the revenues on hit model. We can now proceed to evaluate the model. Let us first assume that the dealer does not expect informational advantage in this RfQ, and there is no other expectation of short-term directionality of the market. We can then use $P_{m,T}-P_{m,t}\sim N(0,\sigma^{2}(T-t))$ and compute the optimal spread:
@@
We can extend this model to the case in which the dealer also has a view on the informational advantage of the client. A simple model that captures the influence of the information asymmetry into the market dynamics is: $\hat{\mu}1_{\text{IA}=1}$. This means that if $\text{IA}=1$, then $P_{m,T}-P_{m,t}\sim N(\hat{\mu}(T-t),\sigma^{2}(T-t))$. The conditioning set to compute revenues in this case does not change, as discussed in the next section. Working out again the utility maximization problem (Eq. 10), the optimal spread has an extra additive correction that compensates for the information asymmetry risk:
@@
To compute this expression, we use the sum and product probability rules to introduce the RfQ status variable ($RS$), and use the fact that $R>0$ only when the dealer trades the RfQ, i.e. $\text{RS}=\text{hit}$:
@@
Factoring the axe matcher into these two probabilities provides a more tractable formulation. Since, in our model, there is no direct causal effect between the client’s demand (captured by the RfQ variable) and the spread $\delta$, assuming the existence of a conditioning set $\mathcal{Z}_{t}$ that blocks all the spurious correlations between these variables —without inducing new ones— allows us to simplify the second probability. A similar argument holds for the RfQ status ($RS$) with respect to the commercial action variable callcall in the first probability. Under these assumptions, the overall axe matcher expression can be rewritten as:
@@
$P(\text{hit}|\text{do}(\delta),\text{RfQ},\text{axe}=1,CF,BF,\mathcal{Z}_{t})\times$
@@
Rewriting the second factor of Eq. 18 in a similar manner, the ACE reads:
$ACE=P(\text{hit}|\text{do}(\delta),\text{RfQ},\text{axe}=1,CF,BF,\mathcal{Z}_{t})\times$
[…]
@@
The first term can be directly linked to the hit probability model, since conditioning on RfQ allows us to remove any influence from the variable axeaxe: at this point, the client’s decision to trade is assumed to depend solely on the pricing, regardless of whether the dealer is axed or not. To estimate this model using historical data, as previously discussed, we must include the volatility $\sigma$ and the RfQ features RFRF—such as the requested volume and the number of competing dealers nn—in the conditioning set. However, since these variables are not known in advance in the context of the axe matcher problem, we must integrate over their conditional probability distribution to properly account for their effect.
@@ 4 Causal analysis of the interventional probabilities
The hit probability model, as discussed in the previous section, is central to the business questions examined in this work. Our objective is to evaluate the causal quantity $P(RS=\text{hit}\mid\text{do}(\delta),\text{RfQ},\mathcal{Z}_{t})$. Specifically, we seek to determine the minimal set of variables in $\mathcal{Z}_{t}$ that enables estimation of this expression from historical (i.e., observational) data. According to Causal Inference theory [5], a sufficient—though not necessary—condition is that the set $\mathcal{Z}_{t}$ satisfies the back-door criterion. Given a causal graph and a pair of variables $X$ and $Y$, a set of variables $\mathcal{Z}$ satisfies this criterion if (i) no variable in $\mathcal{Z}$ is a descendant of $X$, and (ii) $\mathcal{Z}$ blocks all back-door paths from $X$ to YY—that is, all paths that contain an arrow pointing into $X$. If such a set $\mathcal{Z}_{t}$ exists, we can validly equate the interventional and observational expressions:
@@
  • The paths via the RfQ features $RF$. On the one hand, the path $\delta\leftarrow RF\rightarrow R\leftarrow RS$ is blocked since $R$, the revenues, is a collider with respect to any direct or indirect path connecting it to $RS$. By not conditioning on revenues, these paths are blocked. Paths passing through $BF$, CFCF and RfQ are blocked due to the conditioning on those variables. Finally we have paths connecting RfQ features to $RS$ via latent variables like $\delta_{\text{res}}$ and $\delta_{\text{dealer}}$, which are open. Again, to block these spurious influences using observable variables we can include $RF$ in the conditioning set.
@@
$\times P(RF,BF,CF|\text{RfQ})dP(\sigma)$ (24)
@@
$P(R>0|\text{do}(\delta),\text{hit},\text{RfQ},\sigma,RF)=\sum_{\text{IA}=0,1}$ $\;P(R>0|\delta,\text{hit},IA,\sigma,RF)$
$\times P(IA|\delta,\text{hit},\text{RfQ},\sigma,RF)$ (25)
@@ 5 Model specification
So far we have used Causal Inference theory to express business causal interventions in terms of probabilities that can be estimated using historical data. When it comes to actually computing those probabilities, we need to provide a model specification. Here, two main routes can be taken:
@@
  • For the conditional distribution of the spreads quoted by other dealers, we adopt the Skew Exponential Power (SEP) distribution, as proposed in [12], applied to the normalized dealer half-spread ($\delta_{\text{dealer}}/\Delta$). To visually assess the appropriateness of this choice, we leverage a structural property of our model: the conditional independence of dealer quotes given their parent features. Under this assumption, the cover price—defined as $\delta_{\text{cover}}=\max\{\delta_{\text{dealer},1},\dots,\delta_{\text{dealer},n}\}$ for buy-side RfQs (and the minimum for sell-side RfQs)—follows the distribution of the maximum of $n$ independent random variables, each drawn from the same SEP distribution conditional on parent features. Fig 2 displays this distributional fit using the SEP assumption. Notably, despite our dataset differing from the one used in [12], the results offer further evidence of the robustness of the SEP specification. To model the conditional distribution of dealer spreads with respect to their parent nodes—which include bond, client, and RfQ features, as well as volatility—we use a linear model:
@@
  • For the distribution of the price discovery variable $PD$, conditional on its parent variables $BF$ and $CF$, one could, in principle, use Bernoulli distributions. However, this approach risks an exponential increase in the number of parameters if a large number of features is included. To address this issue while preserving interpretability and tractability, we model PDPD—a binary random variable—using a logistic regression model, namely: